# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v2

@@ C.1 Proof of Lemma 1
Since $E=\mathbf{e}d^{\pi\top}$, $E$ is a rank‑one projector onto the span of $\mathbf{e}$. Hence the spectrum of $E$ is $\{1,0,\ldots,0\}$. In addition, we can show $\mathsf{P}^{\pi}$ and $E$ commute by
$\mathsf{P}^{\pi}E=\mathsf{P}^{\pi}(\mathbf{e}d^{\pi})=(\mathbf{e}d^{\pi})=E$ (28)
$E\mathsf{P}^{\pi}=\mathbf{e}(d^{\pi}\mathsf{P}^{\pi})=\mathbf{e}d^{\pi}=E$ (29)
Thus, by the Schur theorem, $\mathsf{P}$ and $E$ are simultaneously upper triangularizable. In a common triangular basis, the diagonals of $\mathsf{P}$ and $E$ list their eigenvalues in descending orders, which are $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{S}\}$ and $\{1,0,\ldots,0\}$ respectively. Thus, in that same basis, $Q^{\pi}=\mathsf{P}^{\pi}-E$ is also triangular, with diagonal entries being $\{\lambda_{1}-1,\lambda_{2}-0,\ldots,\lambda_{S}-0\}$. Since $\lambda_{1}=1$, we have the spectrum of $Q^{\pi}$ is exactly $\{\lambda_{2},\ldots,\lambda_{S},0\}$. Since we already have $|\lambda_{i}|<1$ for all $i\in\{2,\ldots,S\}$, we conclude the proof. ∎
Define $\rho(\cdot)$ to be the spectral radius of a matrix, then Lemma 4 implies that $\rho(Q^{\pi})<1$. Hence by equivalence of norms in $\mathbb{R}^{|\mathcal{S}|}$ it is possible to construct a vector norm $\|\cdot\|_{Q}$ so that the induced operator norm of $Q^{\pi}$ is less than $1$, specifically
$10\leq\rho(Q^{\pi})\leq\bigl{\|}Q^{\pi}\bigr{\|}_{Q}\leq\alpha<$ 1 . (30)
A concrete construction example is to leverage the discrete-Lyapunov equation Horn and Johnson (2012) of solving $M$ on the space of symmetric matrices for any $\rho(Q^{\pi})<\alpha<1$ as follows:
@@
We can express $M$ in the form of Neumann‐series as
@@
$=\sum_{k=0}^{\infty}\alpha^{-2(k+1)}\left((Q^{\pi})^{\top}\right)^{k}(Q^{\pi})^{k}.$ (32)
Since $\rho((Q^{\pi})^{\top}Q^{\pi})=\rho(Q^{\pi})^{2}<\alpha^{2}<1$, the partial sum expression in (C.1) converges absolutely as a geometric‐type series. Also, since each term in (C.1) is positive semi‑definite, and the first term $\alpha^{-2}I$ being positive definite, we can conclude that $M$ being the summation is well-defined and is positive definite. Thus, using the $M$ defined in (31), we can define our desired $\|\cdot\|_{Q}$ as
@@
Based on the above construction of $\|\cdot\|_{Q}$, define the operator $\|\cdot\|_{\mathsf{P}}$ as
@@
Let $\beta=\alpha+\epsilon$, by (30) and (36), we have $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, $\beta\in(0,1)$ and combining $\beta$ with the semi-norm $\|\cdot\|_{\mathsf{P}}$ confirms Lemma 1.
C.2 Proof of Theorem 2

### v3

@@ Appendix A Semi-Norm Contraction Property of the Bellman Operator
Since $\mathbf{e}d^{\pi\top}, EE$ is a rank‑one projector onto the span of $\mathbf{e}$. Hence the spectrum of $E$ is $\{1,0,\ldots,0\}$. In addition, we can show $\mathsf{P}^{\pi}$ and $E$ commute by
$\mathsf{P}^{\pi}E=\mathsf{P}^{\pi}(\mathbf{e}(d^{\pi})^{\top})=(\mathbf{e}(d^{\pi})^{\top})=E$ (28)
$E\mathsf{P}^{\pi}=\mathbf{e}((d^{\pi})^{\top}\mathsf{P}^{\pi})=\mathbf{e}(d^{\pi})^{\top}=E$ (29)
Thus, by the Schur’s theorem, $\mathsf{P}$ and $E$ are simultaneously upper triangularizable. In a common triangular basis, the diagonals of $\mathsf{P}$ and $E$ list their eigenvalues in descending orders, which are $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{S}\}$ and $\{1,0,\ldots,0\}$ respectively. Thus, in that same basis, $Q^{\pi}=\mathsf{P}^{\pi}-E$ is also triangular, with diagonal entries being $\{\lambda_{1}-1,\lambda_{2}-0,\ldots,\lambda_{S}-0\}$. Since $\lambda_{1}=1$, we have the spectrum of $Q^{\pi}$ is exactly $\{\lambda_{2},\ldots,\lambda_{S},0\}$. Since we already have $|\lambda_{i}|<1$ for all $i\in\{2,\ldots,S\}$, we conclude the proof. ∎
Define $\rho(\cdot)$ to be the spectral radius of a matrix, then Lemma A.1 implies that $\rho(Q^{\pi})<1$. Hence by equivalence of norms in $\mathbb{R}^{|\mathcal{S}|}$ it is possible to construct a vector norm $\|\cdot\|_{Q}$ so that the induced operator norm of $Q^{\pi}$ is less than $1$, specifically
$0\leq\rho(Q^{\pi})\leq\bigl\|Q^{\pi}\bigr\|_{Q}\leq\alpha<1.$ (30)
A concrete construction example is to leverage the discrete-Lyapunov equation [19] of solving $M$ on the space of symmetric matrices for any $\rho(Q^{\pi})<\alpha<1$ as follows:
@@
Define $B\coloneqq\alpha^{-1}Q^{\pi}$, then $\rho(B)=\alpha^{-1}\rho(Q^{\pi})<1$. We can express $M$ in the form of Neumann series as
@@
$=\sum_{k=0}^{\infty}\alpha^{-2(k+1)}\left((Q^{\pi})^{\top}\right)^{k}(Q^{\pi})^{k}$
$=\alpha^{-2}\sum_{k=0}^{\infty}\left(B^{\top}\right)^{k}B^{k}.$ (32)
We now show that $M$ is bounded. Write $B=SJS^{-1}$, where $J=\mathrm{diag}\big(J_{m_{1}}(\lambda_{1}),\dots,J_{m_{r}}(\lambda_{r})\big)$ is the Jordan normal form. By the Jordan block power formula [19],
$J_{m}(\lambda)=\lambda I_{m}+N_{m},$ (33)
with $N_{m}$ the nilpotent matrix having ones on the first superdiagonal and $N_{m}^{m}=0$. Then $B^{k}=SJ^{k}S^{-1}$ and $J^{k}=\mathrm{diag}\big(J_{m_{1}}(\lambda_{1})^{k},\dots,J_{m_{r}}(\lambda_{r})^{k}\big)$. For each block and each integer $k\geq m$, by the binomial theorem we have $N_{m}^{m}=0$ and
$J_{m}(\lambda)^{k}=(\lambda I_{m}+N_{m})^{k}=\sum_{j=0}^{m-1}\binom{k}{j}\,\lambda^{\,k-j}\,N_{m}^{\,j}.$ (34)
For $k\geq m$, use $\binom{k}{j}\leq\frac{k^{j}}{j!}$ and factor $|\lambda|^{\,k}$:
$\|J_{m}(\lambda)^{k}\|_{2}\leq|\lambda|^{\,k}\sum_{j=0}^{m-1}\frac{k^{j}}{j!}|\lambda|^{-j}\|N_{m}^{j}\|_{2}\leq c_{m,\lambda}k^{m-1}|\lambda|^{k},$ (35)
where $c_{m,\lambda}\coloneqq\sum_{j=0}^{m-1}\frac{|\lambda|^{-j}}{j!}\|N_{m}^{j}\|_{2}$. Thus, let $s=\max_{i}m_{i}$ be the size of the largest Jordan block of $B$. Since $J^{k}$ is block diagonal,
$\|J^{k}\|_{2}\leq\sum_{i=1}^{r}\|J_{m_{i}}(\lambda_{i})^{k}\|_{2}\leq\Big(\sum_{i=1}^{r}c_{m_{i},\lambda_{i}}\Big)k^{s-1}\Big(\max_{i}|\lambda_{i}|\Big)^{k}=C_{J}k^{s-1}\rho(B)^{k}$ (36)
for all $k\geq s$, where $C_{J}:=\sum_{i=1}^{r}c_{m_{i},\lambda_{i}}$ and $\rho(B)=\max_{i}|\lambda_{i}|$.
Since similarity does not change eigenvalues but may scale norms by the condition number, we can derive that
$\|B^{k}\|_{2}=\|SJ^{k}S^{-1}\|_{2}\leq\|S\|_{2}\|S^{-1}\|_{2}\|J^{k}\|_{2}\leq\kappa(S)\,C_{J}k^{s-1}\rho(B)^{k}\qquad(k\geq s),$
where $\kappa(S):=\|S\|_{2}\|S^{-1}\|_{2}$. By choosing the appropriate constant, the same bound holds for all $k\geq 0$:
$\exists\;C_{B}>0\ \text{such that}\quad\|B^{k}\|_{2}\;\leq\;C_{B}\,k^{s-1}\,\rho(B)^{\,k}\quad(k\geq 0).$
In spectral norm, this implies
$\|(B^{\top})^{k}B^{k}\|_{2}=\|(B^{k})^{\top}B^{k}\|_{2}=\sigma_{\max}(B^{k})^{2}=\|B^{k}\|_{2}^{2}\leq C_{B}^{2}k^{2(s-1)}\rho(B)^{2k}.$ (37)
Thus, the scalar series $\sum_{k=0}^{\infty}k^{2(s-1)}\,\rho(B)^{\,2k}$ is in the form of polynomial times geometric with ratio less than $1$, which converges, and the partial sum expression in (A.1) converges absolutely as a geometric‐type series.
Also, since each term in (A.1) is positive semi‑definite, and the first term $\alpha^{-2}I$ being positive definite, we can conclude that $M$ being the summation is well-defined and is positive definite. Thus, using the positive definite $M$ defined in (31), we can define our desired norm $\|\cdot\|_{Q}$ as
@@
Based on the above construction of the norm $\|\cdot\|_{Q}$, define the operator $\|\cdot\|_{\mathsf{P}}$ as
@@
Let $\beta=\alpha+\epsilon$, by (30) and (41), we have $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, $\beta\in(0,1)$ and combining $\beta$ with the semi-norm $\|\cdot\|_{\mathsf{P}}$ confirms Lemma 4.1.
A.2 Proof of Theorem 4.2